Find the slope and the y-intercept of the equation y = 9 + 3/7x

Answer by nerdybill(7384)

Find the slope and the y-intercept of the equation y = 9 + 3/7x
.
The "slope-intercept" form of a linear equation is:
y = mx + b
where
m is the slope
b is the y-intercept
.
We can rewrite your equation:
y = 9 + 3/7x
as
y = 3/7x + 9
.
Then, from inspection, we see that
m = 3/7
b = 9
.
So,
Slope is 3/7
y-intercept is at (0,9)